### v3

# ATLAS: Adaptive Topology-based Learning at Scale for Homophilic and Heterophilic Graphs

Abstract
We present ATLAS (Adaptive Topology-based Learning at Scale for Homophilic and Heterophilic Graphs), a novel graph learning algorithm that addresses two important challenges in graph neural networks (GNNs). First, the accuracy of GNNs degrades when the graph is heterophilic. Second, the iterative feature aggregation limits the scalability of GNNs on large graphs. We address these challenges by extracting topological information about the graph communities at different levels of refinement, concatenating the community assignments to the feature vector, and applying multilayer perceptrons (MLPs) on this new feature vector. We thus inherently obtain the topological data about the nodes and their neighbors without invoking aggregation. Because MLPs are typically more scalable than GNNs, our approach applies to large graphs, without the need for sampling.
Our results, on a wide set of graphs, show that ATLAS has comparable accuracy to baseline methods, with accuracy being as high as $20$ percentage points over GCN for heterophilic graphs with negative structural bias and $11$ percentage points over MLP for homophilic graphs. Furthermore, we show how multi-resolution community features systematically modulate performance in both homophilic and heterophilic settings, opening a principled path toward explainable graph learning.
1 Introduction
Node classification, a fundamental problem in graph learning, involves identifying labels of nodes in a graph and has wide applications in many domains including social networks, citation networks, recommendation systems, knowledge graphs and bioinformatics (Khemani, 2024; Wu et al., 2019b; Zhou et al., 2021). Accurate classification requires two complementary pieces of information–(i) the features at each node, and (ii) the connections between the node and its neighbors. Neural network methods such as Multi-Layer Perceptrons (MLPs) are fast but do not include information about the connections. Graph Neural Networks (GNNs) address this problem by aggregating the features between neighboring nodes, but the process is expensive and difficult to scale to large graphs. To improve scalability, sampling-based methods construct mini-batches by sampling neighborhoods or subgraphs (e.g., GraphSAGE, ClusterGCN, GraphSAINT) (Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2020); however, this introduces an additional granularity choice—the fan-out or subgraph/cluster size—that trades off computational cost against how much structural context is captured per update. Although the graph structure can be represented as feature vectors using different node embedding techniques (Perozzi et al., 2014; Grover and Leskovec, 2016; Tang et al., 2015), or through the use of community-aware representations/features derived from community detection (Sun et al., 2019; Kamiński et al., 2024), the issue remains as to how many hops of neighbors should be considered and how fine-grained the communities (or sampled subgraphs) should be. Larger hops, larger sampled subgraphs, or coarse-grained communities can lead to information smoothing, while smaller hops, smaller sampled subgraphs, or fine-grained communities can lead to information loss. Further, the hypothesis that aggregating features of neighbors can improve accuracy of node classification is only true for homophilic networks (where nodes of similar classes are connected). In heterophilic networks, where the connection between nodes need not imply similarity of class, this strategy leads to lower accuracy. Based on these observations, we posit that matching structural information (i.e., size of hops, sampled subgraphs/clusters, or communities) with how well it aligns with the classification task is necessary for producing accurate results.
1.1 Related Work
Graph Neural Networks (GNNs) have become a core tool for learning on graphs (Kipf and Welling, 2017; Hamilton et al., 2017). Most algorithms follow a message-passing paradigm, aggregating transformed neighbor features into topology-aware embeddings, which implicitly assumes homophily (Wu et al., 2019a). The same bias can blur informative distinctions on weakly homophilous or heterophilous graphs (Zhu et al., 2020; Platonov et al., 2023a).
Scaling GNNs on large graphs. Scaling GNNs on large graphs is challenging due to memory and aggregation costs. Sampling-based methods approximate full-batch propagation using node-, layer-, or subgraph-level sampling (GraphSAGE, FastGCN, Cluster-GCN, GraphSAINT, LABOR) (Hamilton et al., 2017; Chen et al., 2018b; Chiang et al., 2019; Zeng et al., 2020; Balın and Çatalyürek, 2023), but introduce stochasticity that affects convergence and reproducibility (Chen et al., 2018b; Zou et al., 2019). Decoupled models instead precompute feature diffusion and train MLPs on fixed graph-derived features, enabling i.i.d. node mini-batching and fast inference (SGC, SIGN, SAGN, GAMLP, SCARA, LD2) (Wu et al., 2019a; Rossi et al., 2020; Sun et al., 2021; Chien et al., 2022; Liao et al., 2022; 2023).
Learning on non-homophilous graphs. For non-homophilous graphs, one line of work preserves self-features while carefully injecting neighborhood information (H2GCN, GloGNN) (Zhu et al., 2020; Li et al., 2022), or reweights neighbors to downweight harmful edges (GPR-GNN, FAGCN) (Chien et al., 2021; Bo et al., 2021). Others exploit higher-order propagation or spectral filters to capture both homophilic and heterophilic signals (MixHop, JacobiConv, BernNet, GBK-GNN) (Abu-El-Haija et al., 2019; Wang and Zhang, 2022; He et al., 2021; Du et al., 2022). See Zheng et al. (2022); Luan et al. (2024b) for broader surveys.
Community-aware node embeddings. Several works use community structure as an explicit representation for downstream prediction. Sun et al. (2019) propose vGraph, a generative model that jointly infers discrete communities and continuous node embeddings by reconstructing edges, so that community assignments act as latent variables guiding representation learning. Closer to our setting, Kamiński et al. (2024) construct community-aware node features (e.g., counts and statistics over community memberships in a node’s ego-network) and feed them into standard classifiers, showing that purely community-derived signals can already yield strong performance on node-level tasks.
Graph–task alignment and community structure. A related line of work asks when a graph’s communities are informative for the labels and how this alignment controls the benefit of message passing. Hussain et al. (2021) vary homophily and community structure in real graphs and define a measure of label–community correlation, showing that GNN gains are largest when labels follow communities and can vanish when they do not. This links the classical “cluster assumption” in semi-supervised learning (Chapelle et al., 2006) with recent analyses of graph–task and NTK–graph alignment in GNN training dynamics (Yang et al., 2024), and motivates methods that treat communities as task-relevant structural signals.
Community-guided graph rewiring. Building on modularity-based detection (Newman, 2006; Blondel et al., 2008), ComMa and ComFy (Rubio-Madrigal et al., 2025) use community structure and feature similarity to rewire intra- and inter-community edges, improving label–community alignment and GNN accuracy on both homophilic and heterophilic graphs.
Unlike community-aware GNNs and rewiring methods, ATLAS treats multi-resolution community assignments as features for a simple MLP, remaining propagation-free while still leveraging community structure.
1.2 Our Contribution
Most of the current research either focuses primarily on homophilic graphs, or the processes to address the heterophilic graphs require expensive operations, such as signal identification/modification, rewiring or spectral gap maximization. These methods cannot efficiently scale to large graphs. Our primary contribution is to develop Adaptive Topology -based Learning at Scale (ATLAS)  [^1] Apart from the acronym, the name ATLAS is to convey our method can handle different degrees of homophily, similar to how an atlas encompasses all different countries., a novel graph learning algorithm that can produce high-accuracy results for both homophilic and heterophilic graphs. ATLAS is based on a simple but powerful technique of refining communities in networks to match the degree of homophily.
Rationale. Our algorithm is based on quantifying homophily through the lens of normalized mutual information (NMI). Given two partitions of the same set of elements NMI measures how well the partitions correspond to each other. If we consider one partition as the communities in the graph, and the other partition as labels, then NMI provides a measure for the degree of homophily in the graph. ATLAS focuses on refining/coarsening communities to identify the region of highest NMI—which will correspond to the highest accuracy. Figure 1 provides an overview of ATLAS.
Our specific contributions are:
  1. 1.
    Theory. We provide a theoretical analysis of how refining communities changes in NMI (Section 2).
  2. 2.
    Algorithm. Based on this mathematical understanding, we develop our algorithm ATLAS (Section 3).
  3. 3.
    Experiments. Provide extensive empirical evaluations by comparing ATLAS across a mix of 13 (8 medium size and 5 large) homophilic and heterophilic graphs, and 14 (9) GNN/MLP-based algorithms for medium sized (large) graphs (Section  4).
  4. 4.
    Analysis. We study how performance varies with community refinement and show consistent trends across “high/low/negative structural-bias” regimes, offering an interpretable view of when topology helps (Section  5).
  5. 5.
    Bridging Frameworks. Unlike prior MLP-based models designed primarily for heterophilic graphs, ATLAS effectively supports both homophilic and heterophilic settings, thereby minimizing the accuracy gap traditionally observed between MLPs and GNNs.
Figure 1: Overview of the community-augmented feature learning pipeline. Community assignments at multiple resolutions are one-hot encoded, projected, concatenated with node features, and input to an MLP for classification.
2 Theoretical Analysis
We mathematically show how refining communities leads to changes in NMI. We define some terms that will help us in the analysis. The proofs of the theorems are given in the appendix.
Let $N$ be the set of nodes. Let $P=\{P_{1},\dots,P_{k}\}$ be a partition of $N$; i.e.
$P_{i}\neq\varnothing,\quad P_{i}\cap P_{j}=\varnothing\ (i\neq j),\quad\text{and}\quad\bigcup_{i=1}^{k}P_{i}=N.$
Let $S=\{S_{1},\dots,S_{m}\}$ be another partition of $N$. We say $S$ is a refinement of $P$ (denoted as $S\preceq P$) iff every block of $S$ is contained in some block of $P$. Formally:
$S\preceq P\quad\Longleftrightarrow\quad\forall\,S_{j}\in S\ \exists\,P_{i}\in P\ \text{such that}\ S_{j}\subseteq P_{i}.$
Normalized mutual information (NMI) is a popular measure to quantify alignment between two partitions. Given two partitions $P$ and $Q$, over a set of $N$ elements and $n_{ij}=|P_{i}\cap Q_{j}|,n_{i}=|P_{i}|,n_{j}=|Q_{j}|$ their normalized mutual information is given as;
$\text{NMI}(P,Q)\;=\;\frac{2I(P;Q)}{H(P)+H(Q)}$
$I(P;Q)\;=\;\sum_{i=1}^{k}\sum_{j=1}^{m}\frac{n_{ij}}{N}\log\left(\frac{N\,n_{ij}}{n_{i}\,n_{j}}\right)$ is the mutual information between partitions $P$ and $Q$. This quantity measures how much information is shared between the partitions $P$ and $Q$. The higher the value, the better the alignment between the partitions. $H(P)=-\sum_{i=1}^{k}\frac{n_{i}}{N}\log\left(\frac{n_{i}}{N}\right),$ is the entropy of partition $P$. $H(Q)$ is defined similarly. The entropy measures the distribution of points in each partition. Low entropy means data is concentrated in few clusters, and is indicative of good clustering.
The value of NMI ranges from 1 (indicating complete alignment between partitions) to close to 0 (indicating complete mismatch between partitions). NMI is high if the partitions are well matched ($I(P,Q)$ is high), and entropy is low ($H(P)$, $H(Q)$ is low).
Figure 2: Effect of refinement on NMI. Initially when clusters have mixed items, NMI is low. The first refinement matches the items and clusters, increasing the NMI. Further refinement does not improve the alignment (mutual information), but increases the spread (entropy), thus decreasing NMI.
Lemma 1 (Refinement does not decrease mutual information).
Let $L$ be labels and $C$ a community partition. Let $C^{\prime}$ be a refinement of $C$, i.e., $C^{\prime}\preceq C$. Then $I(L;C^{\prime})\geq I(L;C)$
Lemma 2 (Refinement does not decrease entropy).
Let $C$ a community partition. Let $C^{\prime}$ be a refinement of $C$, i.e., $C^{\prime}\preceq C$. Then $H(C^{\prime})\geq H(C)$
Based on Lemma 1 and Lemma 2 we see that while refinement improves the mutual information leading to better alignment, it also increases the entropy leading to more noise or uncertainty. The condition at which NMI will increase is given by Theorem 1.
Theorem 1 (NMI Refinement Condition).
Let $L$ be labels; $C$ a community partition. Let $C^{\prime}$ be a refinement of $C$, i.e., $C^{\prime}\preceq C$. Then $\text{NMI}(C^{\prime};L)>\text{NMI}(C;L)\quad\text{if and only if}\quad\frac{\Delta I}{\Delta H}>\frac{\text{NMI}(C;L)}{2}$; where $\Delta I=I(C^{\prime};L)-I(C;L)$ and $\Delta H=H(C^{\prime};L)-H(C;L)$
Theorem 1 states that a partition refinement improves the normalized mutual information with respect to labels if and only if the mutual information gain per unit of entropy increase exceeds half the original normalized mutual information value.
3 Methodology
The theorems in Section 2 are based on idealized conditions, where refined communities are perfect subsets of the original communities. In practice, refinement in communities is approximated by running a modularity-based community detection algorithm at multiple resolution values. Although higher resolution leads to smaller communities, due to the inherent non-determinism of community detection methods, the smaller communities may not be exact subsets.
Preprocessing. Optimizing modularity is a popular method for community detection. Modularity, $Q$, measures the strength of connections between nodes in a community as compared to a null model with randomly placed edges. Communities in networks are often hierarchical, so we treat the resolution parameter $\gamma$ as the hierarchy/refinement level (larger $\gamma$ yields finer-grained communities); Appendix 8.1 summarizes the community terminology and formal definitions used below. We start from two initial resolutions ($\gamma=0.5$ and $\gamma=1.0$) and set three hyperparameters: a modularity gap threshold $\Delta_{\max}$, a minimum modularity $Q_{\min}$, and a small target-drop range $[a,b]$. Let $\gamma_{1}$ and $\gamma_{2}$ be two consecutive resolution parameters, with community sets $\mathbf{c}^{(\gamma_{1})}$ and $\mathbf{c}^{(\gamma_{2})}$ and modularities $Q^{(\gamma_{1})}$ and $Q^{(\gamma_{2})}$; we define the modularity gap as $\Delta Q=\lvert Q^{(\gamma_{2})}-Q^{(\gamma_{1})}\rvert$. At each iteration, we sort the tested resolutions and examine consecutive pairs. If the gap between a pair exceeds $\Delta_{\max}$, we find their midpoint (interpolation). Otherwise, we extrapolate beyond the current maximum by estimating the local slope of modularity with respect to the resolution and taking a small forward step expected to reduce modularity by a random amount drawn from the drop range. Once the new $\gamma$ is obtained, we compute the communities at that value. The loop stops when the latest modularity falls below $Q_{\min}$ or no new resolution is produced. The procedure returns the retained resolutions and their corresponding community assignments, which we view as multi-resolution community features—a structural graph signal over the nodes—that are later encoded and concatenated with the original node features in the feature-augmentation step (see Algorithm 1 in the appendix).
Figure 3: Illustration of the Adaptive Resolution Search Process. The resolution limits, $\gamma_{0}<\gamma_{2}<\gamma_{1}<\gamma_{3}<\gamma_{4}$, and the communities $C^{\gamma_{0}}\preceq C^{\gamma_{2}}\preceq C^{\gamma_{1}}\preceq C^{\gamma_{3}}\preceq C^{\gamma_{4}}$ capture structural bias for different granularities from the graph.
Feature Augmentation. For a resolution $\gamma\in\Gamma$, community detection on $G$ yields node assignments $\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}^{n}$. We one-hot encode them as $\mathbf{H}^{(\gamma)}$ (equation 2) and project to a dense space using a trainable $\mathbf{W}^{(\gamma)}$ (equation 3). Concatenating embeddings across resolutions forms $\mathbf{E}$ (equation 4), which is concatenated with node features $\mathbf{X}$ to obtain $\mathbf{Z}_{\text{ATLAS}}$. An MLP $f_{\theta}$ outputs logits, and a task-dependent map $\phi$ (e.g., softmax or elementwise sigmoid) produces probabilities $\hat{\mathbf{Y}}$ (equation 6).
$\mathbf{c}^{(\gamma)}$ $=\mathrm{DetectCommunity}(G,\gamma),\quad\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}^{n}$ (1)
$\mathbf{H}^{(\gamma)}$ $=\mathrm{OneHot}(\mathbf{c}^{(\gamma)})\in\{0,1\}^{n\times k_{\gamma}}$ (2)
$\mathbf{E}^{(\gamma)}$ $=\mathbf{H}^{(\gamma)}\mathbf{W}^{(\gamma)},\quad\mathbf{W}^{(\gamma)}\in\mathbb{R}^{k_{\gamma}\times d_{c}}$ (3)
$\mathbf{E}$ $=\big\|_{t=1}^{T}\mathbf{E}^{(\gamma_{t})}\in\mathbb{R}^{n\times(Td_{c})}$ (4)
$\mathbf{Z}_{\text{ATLAS}}$ $=[\,\mathbf{X}\ \|\ \mathbf{E}\,]\in\mathbb{R}^{n\times(D+Td_{c})}$ (5)
$\hat{\mathbf{Y}}$ $=\phi\!\big(f_{\theta}(\mathbf{Z}_{\text{ATLAS}})\big)\in[0,1]^{n\times C}.$ (6)
Local neighbor effects.
Multi-resolution community embeddings capture informative meso-scale structure, but may under-represent strictly local signals driven by immediate neighbors (e.g., mixed labels within communities or short-range dependencies), as well as fine-grained attribute context. Without changing the ATLAS backbone, we therefore evaluate two optional neighborhood-based augmentations computed from the original adjacency $\mathbf{A}$. ATLAS-NF aggregates 1-hop neighbor features, $\mathbf{X}_{\mathrm{NF}}=\mathbf{A}\mathbf{X}$. ATLAS-LPF seeds training labels only (one-/multi-hot) and diffuses them using degree-normalized propagation $\tilde{\mathbf{A}}=\mathrm{RowNorm}(\mathbf{A})$ to obtain a soft neighborhood label prior; we concatenate NF/LPF with $\mathbf{Z}_{\text{ATLAS}}$.
ATLAS-NF: $\mathbf{X}_{\mathrm{NF}}=\mathbf{A}\mathbf{X},\qquad\mathbf{Z}_{\mathrm{ATLAS\text{-}NF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{X}_{\mathrm{NF}}\,].$ (7)
ATLAS-LPF: $\mathbf{Y}^{(0)}=(\mathbf{m}_{\mathrm{tr}}\mathbf{1}_{C}^{\top})\odot\mathbf{Y},\qquad\mathbf{Y}^{(h)}=\tilde{\mathbf{A}}\mathbf{Y}^{(h-1)},\ h=1,\dots,H,$
$\mathbf{Y}_{\mathrm{LPF}}=\big\|_{h=1}^{H}\mathbf{Y}^{(h)}\in\mathbb{R}^{n\times(HC)},\qquad\mathbf{Z}_{\mathrm{ATLAS\text{-}LPF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{Y}_{\mathrm{LPF}}\,].$
$\mathbf{Z}_{\mathrm{ATLAS\text{-}NF\text{-}LPF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{X}_{\mathrm{NF}}\|\mathbf{Y}_{\mathrm{LPF}}\,].$
Complexity Analysis. We compare computational and memory complexities of representative scalable GNN frameworks with our approach in Table 3. ATLAS performs Louvain clustering in $O(T\|A\|_{0})$ in the preprocessing step, keeps a single augmented feature buffer, and trains with per-epoch time $O\!\big(L_{ff}N(D{+}Td_{c})^{2}\big)$ and memory $O\!\big(bL_{ff}(D{+}Td_{c})\big)$, enabling simple i.i.d. node mini-batching without neighborhood expansion or graph-dependent batching heuristics. ATLAS performs adjacency-free inference: with fixed augmented features of dimension $D{+}Td_{c}$, prediction is a forward pass with complexity $O\!\big(N(D{+}Td_{c})^{2}\big)$.
4 Empirical Evaluation
In this section, we provide the empirical results comparing ATLAS with other graph learning methods. Our experiments focus on answering the following research questions:
Q1. How accurate is ATLAS compared to baseline methods over graphs with different degrees of homophily?
Q2. How well can ATLAS scale to large graphs, while maintaining high accuracy?
Datasets. We use 8 medium graphs (Cora, PubMed, Tolokers, Squirrel-Filtered, Chameleon-Filtered, Amazon-Ratings, Actor, Roman-Empire) and 5 large graphs (Flickr, Reddit, Yelp, Amazon-Products, OGBN-Products). Complete statistics of datasets are given in Appendix Tables 8 and 9.
Baselines. We group baselines by modeling regime and map them to the research questions.
Q1 (homophily–heterophily regime). Homophilic: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018). Heterophily-oriented: H2GCN (Zhu et al., 2020), LinkX (Lim et al., 2021), GPR-GNN (Chien et al., 2021), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang and Zhang, 2022), ACM-GCN (Luan et al., 2022), BernNet (He et al., 2021).
Q2 (scalability). Decoupled: SGC (Wu et al., 2019a), SIGN (Rossi et al., 2020), SAGN (Sun et al., 2021), GAMLP (Chien et al., 2022). Sampling-based: GraphSAGE (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), GraphSAINT (Zeng et al., 2020). Descriptions of these methods are provided in the Appendix.
We use an $L$-layer MLP with hidden width $d_{\mathrm{hid}}$ and dropout rate $p$. Each of the first L1L{-}1 layers applies Linear (with bias) $\rightarrow$ LayerNorm $\rightarrow$ GELU $\rightarrow$ Dropout. The final layer is a Linear classifier to $C$ classes.
4.1 Q1: Accuracy Across Homophily Regimes
Table 1 summarizes results across three homophily regimes (high, low, negative). On high-homophily graphs, ATLAS improves over a feature-only MLP (e.g., +11.65 on Cora) and remains competitive with strong GNN baselines; ATLAS-NF is best on Tolokers (83.02). On low-homophily graphs, ATLAS achieves the top result on Chameleon-Filtered (42.76), while ATLAS-LPF is best on Amazon-Ratings (53.18), indicating that a local label prior can complement meso-scale community signals. On negative-homophily graphs, ATLAS consistently outperforms GCN (e.g., +20.54 on Roman-Empire) and local augmentations further help on feature-driven cases: ATLAS-NF reaches 78.10 and ATLAS-LPF-NF attains the best overall accuracy on Roman-Empire (80.80). Overall, ATLAS is robust across regimes, with NF/LPF improving cases where strictly local attribute or neighborhood-label information is critical.
Table 1: Eight-benchmark comparison across homophily regimes. Baseline heterophily-oriented model results are from Platonov et al. (2023b); Luan et al. (2024a). Bottom rows report ATLAS improvements over baselines (absolute percentage points). Cells highlighted in yellow indicate the best score for each dataset.
High structural bias Low structural bias Negative structural bias
Model Cora $h_{e}=0.810$ Tolokers $h_{e}=0.595$ PubMed $h_{e}=0.802$ Chameleon-Filtered $h_{e}=0.236$ Amazon-Ratings $h_{e}=0.380$ Actor $h_{e}=0.216$ Squirrel-Filtered $h_{e}=0.207$ Roman-Empire $h_{e}=0.047$
MLP(2L) 75.44 $\pm$ 1.97 72.97 $\pm$ 0.90 87.25 $\pm$ 0.41 36.00 $\pm$ 4.69 39.83 $\pm$ 0.48 34.96 $\pm$ 0.71 34.29 $\pm$ 3.34 65.58 $\pm$ 0.34
GCN 87.01 $\pm$ 1.04 74.93 $\pm$ 1.32 86.71 $\pm$ 0.42 37.11 $\pm$ 3.04 42.78 $\pm$ 0.14 28.49 $\pm$ 0.91 32.70 $\pm$ 1.73 45.68 $\pm$ 0.38
SAGE 87.50 $\pm$ 0.87 80.95 $\pm$ 0.92 88.42 $\pm$ 0.55 38.83 $\pm$ 4.26 44.67 $\pm$ 0.51 34.08 $\pm$ 1.07 33.32 $\pm$ 1.75 76.21 $\pm$ 0.65
GAT 87.74 $\pm$ 0.88 75.31 $\pm$ 1.35 86.18 $\pm$ 0.64 37.18 $\pm$ 3.44 43.25 $\pm$ 0.85 29.11 $\pm$ 1.23 32.61 $\pm$ 2.06 47.16 $\pm$ 0.66
H2GCN 87.52 $\pm$ 0.61 73.35 $\pm$ 1.01 87.78 $\pm$ 0.28 26.75 $\pm$ 3.64 36.47 $\pm$ 0.23 38.85 $\pm$ 1.17 35.10 $\pm$ 1.15 60.11 $\pm$ 0.52
LinkX 82.62 $\pm$ 1.44 81.15 $\pm$ 1.23 88.12 $\pm$ 0.47 42.34 $\pm$ 4.13 52.66 $\pm$ 0.64 35.64 $\pm$ 1.36 40.10 $\pm$ 2.21 56.15 $\pm$ 0.93
GPR-GNN 79.51 $\pm$ 0.36 72.94 $\pm$ 0.97 85.07 $\pm$ 0.09 39.93 $\pm$ 3.30 44.88 $\pm$ 0.34 39.30 $\pm$ 0.27 38.95 $\pm$ 1.99 64.85 $\pm$ 0.27
FSGNN 87.51 $\pm$ 1.21 82.76 $\pm$ 0.61 90.11 $\pm$ 0.43 40.61 $\pm$ 2.97 52.74 $\pm$ 0.83 37.65 $\pm$ 0.79 35.92 $\pm$ 1.32 79.92 $\pm$ 0.56
GloGNN 87.67 $\pm$ 1.16 73.39 $\pm$ 1.17 90.32 $\pm$ 0.54 25.90 $\pm$ 3.58 36.89 $\pm$ 0.14 39.65 $\pm$ 1.03 35.11 $\pm$ 1.24 59.63 $\pm$ 0.69
FAGCN 88.85 $\pm$ 1.36 77.75 $\pm$ 1.05 89.98 $\pm$ 0.54 41.90 $\pm$ 2.72 44.12 $\pm$ 0.30 31.59 $\pm$ 1.37 41.08 $\pm$ 2.27 65.22 $\pm$ 0.56
GBK-GNN 87.09 $\pm$ 1.52 81.01 $\pm$ 0.67 88.88 $\pm$ 0.44 39.61 $\pm$ 2.60 45.98 $\pm$ 0.71 38.47 $\pm$ 1.53 35.51 $\pm$ 1.65 74.57 $\pm$ 0.47
JacobiConv 89.61 $\pm$ 0.96 68.66 $\pm$ 0.65 89.99 $\pm$ 0.39 39.00 $\pm$ 4.20 43.55 $\pm$ 0.48 37.48 $\pm$ 0.76 29.71 $\pm$ 1.66 71.14 $\pm$ 0.42
BernNet 88.52 $\pm$ 0.95 77.00 $\pm$ 0.65 88.48 $\pm$ 0.41 40.90 $\pm$ 4.06 44.64 $\pm$ 0.56 41.79 $\pm$ 1.01 41.18 $\pm$ 1.77 65.56 $\pm$ 1.34
ACM-GCN 89.75 $\pm$ 1.16 74.95 $\pm$ 1.16 90.96 $\pm$ 0.62 42.73 $\pm$ 3.59 52.49 $\pm$ 0.24 41.86 $\pm$ 1.48 42.35 $\pm$ 1.97 71.89 $\pm$ 0.61
ATLAS 88.37 $\pm$ 0.63 82.19 $\pm$ 0.73 88.85 $\pm$ 0.48 42.76 $\pm$ 3.47 53.15 $\pm$ 0.61 38.48 $\pm$ 0.93 40.82 $\pm$ 2.12 66.22 $\pm$ 0.53
ATLAS-NF 89.20 $\pm$ 0.45 83.02 $\pm$ 0.74 88.76 $\pm$ 0.36 40.02 $\pm$ 2.79 52.30 $\pm$ 0.64 34.26 $\pm$ 0.97 36.98 $\pm$ 2.37 78.10 $\pm$ 0.59
ATLAS-LPF 88.25 $\pm$ 0.46 82.06 $\pm$ 0.61 88.80 $\pm$ 0.12 41.54 $\pm$ 3.44 53.18 $\pm$ 0.30 38.93 $\pm$ 0.65 40.08 $\pm$ 2.19 75.41 $\pm$ 2.20
ATLAS-LPF-NF 89.06 $\pm$ 0.39 82.79 $\pm$ 0.78 88.58 $\pm$ 0.29 40.48 $\pm$ 3.53 52.28 $\pm$ 0.25 36.26 $\pm$ 0.81 38.20 $\pm$ 2.35 80.80 $\pm$ 0.84
ATLAS–MLP +12.93 +9.22 +1.60 +6.76 +13.32 +3.52 +6.53 +0.64
ATLAS–GCN +1.36 +7.26 +2.14 +5.65 +10.37 +9.99 +8.12 +20.54
ATLAS–Average +2.20 +5.97 +0.40 +4.99 +8.51 +2.13 +4.54 +1.67
Table 2: Large-graph performance. Baselines from Zeng et al. (2020); Hu et al. (2020). Bottom rows report ATLAS improvements over MLP and over GCN (absolute percentage points). Cells highlighted in yellow indicate the best score for each dataset.
High structural bias Low structural bias Negative structural bias
Method Reddit $h_{e}{=}0.756$ ogbn-products $h_{e}{=}0.808$ Flickr $h_{e}{=}0.319$ Yelp $h_{e}{=}0.809$ AmazonProducts $h_{e}{=}0.116$
MLP 74.35 $\pm$ 0.16 61.06 $\pm$ 0.08 47.17 $\pm$ 0.11 65.46 $\pm$ 0.11 82.04 $\pm$ 0.02
GCN 93.30 $\pm$ 0.01 75.64 $\pm$ 0.21 49.20 $\pm$ 0.30 37.80 $\pm$ 0.10 28.10 $\pm$ 0.50
GraphSAGE 95.30 $\pm$ 0.10 80.61 $\pm$ 0.16 50.10 $\pm$ 1.30 63.40 $\pm$ 0.60 75.80 $\pm$ 0.20
ClusterGCN 95.40 $\pm$ 0.10 78.62 $\pm$ 0.61 48.10 $\pm$ 0.50 60.90 $\pm$ 0.50 75.90 $\pm$ 0.80
GraphSAINT 96.60 $\pm$ 0.10 75.36 $\pm$ 0.34 51.10 $\pm$ 0.10 65.30 $\pm$ 0.30 81.50 $\pm$ 0.10
SGC 93.51 $\pm$ 0.04 67.48 $\pm$ 0.11 50.35 $\pm$ 0.05 23.56 $\pm$ 0.02 22.62 $\pm$ 0.28
SIGN 95.95 $\pm$ 0.02 80.52 $\pm$ 0.16 51.60 $\pm$ 0.11 57.98 $\pm$ 0.12 74.24 $\pm$ 0.02
SAGN 96.48 $\pm$ 0.03 81.21 $\pm$ 0.07 50.07 $\pm$ 0.11 61.55 $\pm$ 0.40 76.82 $\pm$ 1.15
GAMLP 96.73 $\pm$ 0.03 83.76 $\pm$ 0.19 52.58 $\pm$ 0.12 57.84 $\pm$ 1.54 75.99 $\pm$ 0.26
ATLAS 95.74 $\pm$ 0.04 79.98 $\pm$ 0.14 51.56 $\pm$ 0.14 65.46 $\pm$ 0.11 82.04 $\pm$ 0.02
ATLAS-NF 94.10 $\pm$ 0.07 75.07 $\pm$ 0.30 52.72 $\pm$ 0.11 57.40 $\pm$ 0.21 77.83 $\pm$ 0.10
ATLAS-LPF 96.74 $\pm$ 0.03 80.52 $\pm$ 0.13 51.92 $\pm$ 0.14 60.20 $\pm$ 1.34 56.64 $\pm$ 0.28
ATLAS-LPF-NF 94.47 $\pm$ 0.10 75.14 $\pm$ 0.30 52.82 $\pm$ 0.16 51.92 $\pm$ 2.07 53.70 $\pm$ 0.48
ATLAS–MLP +21.39 +18.92 +4.39 +0.00 +0.00
ATLAS–GCN +2.44 +4.34 +2.36 +27.66 +53.94
ATLAS–Average +2.67 +3.95 +1.53 +10.59 +16.15
Table 3: Complexity comparison. NN = #nodes, $\|A\|_{0}$ = #edges, $D$ = feature dim, $L$ = #message-passing layers, $L_{ff}$ = #feed-forward layers, $b$ = batch size, $r$ = sampled neighbors (or filter size), $K$ = #precomputed hop propagations (max hop order for SAGN/GAMLP), $k$ = #subgraph samples used in GraphSAINT preprocessing, $T$ = #resolutions, $d_{c}$ = community-embedding dim.
Method Preprocessing Per-epoch Train Time Memory
GCN (full-batch) $O(L\,\|A\|_{0}D+LND^{2})$ $O(LND+LD^{2})$
ClusterGCN $O(\|A\|_{0})$ $O(L\,\|A\|_{0}D+LND^{2})$ $O(bLD+LD^{2})$
GraphSAINT $O(kN)$ $O(L\,\|A\|_{0}D+LND^{2})$ $O(bLD)$
SAGN $O(K\|A\|_{0}D)$ $O(L_{ff}N(KD)^{2})$ $O(bL_{ff}KD)$
GAMLP $O(K\|A\|_{0}D)$ $O(L_{ff}N(KD)^{2})$ $O(bL_{ff}KD)$
ATLAS $O(T\|A\|_{0})$ $O(L_{ff}N(D+Td_{c})^{2})$ $O(bL_{ff}(D+Td_{c}))$
Figure 4: The convergence landscape of ATLAS.
Table 4: Computation time breakdown (seconds) on OGBN-Products.
Model Preprocessing Time Per-epoch Train Time Inference Time
GCN $2.0395\pm 0.0006$ $0.9220\pm 0.0010$
ClusterGCN $168.754\pm 1.777$ $4.017\pm 0.164$ $82.837\pm 0.622$
GraphSAINT $3.770\pm 0.159$ (per epoch) $0.751\pm 0.046$ $66.445\pm 0.517$
SAGN $4.8462\pm 0.0415$ $0.8447\pm 0.0225$ $0.2564\pm 0.0001$
GAMLP $4.8227\pm 0.0871$ $0.7976\pm 0.0099$ $0.2495\pm 0.0001$
ATLAS $391.894\pm 14.387$ $0.181\pm 0.0058$ $0.526\pm 0.0038$
4.2 Q2: Efficiency and Scalability on Large Graphs
Accuracy.
Table 2 shows that ATLAS remains competitive at million-node scale while maintaining our propagation-free inference design: prediction is performed by an MLP over topology-derived features (multi-resolution community summaries), optionally augmented with precomputed local signals. On high structural-bias graphs, ATLAS attains $95.74$ (Reddit) and $79.98$ (ogbn-products), and ATLAS-LPF further improves performance to $96.74$ and $80.52$ by injecting a degree-normalized local label prior from training labels. On Flickr (low structural bias), local attribute context is more beneficial: ATLAS-NF improves over ATLAS ($52.72$ vs. $51.56$) and ATLAS-LPF-NF performs best among ATLAS variants ($52.82$). In contrast, on Yelp (negative structural bias), neighborhood aggregation is less label-aligned and NF/LPF reduce accuracy (ATLAS: $65.46$ vs. $57.40$ for ATLAS-NF and $60.20$ for ATLAS-LPF). Overall, these results support the structural-bias perspective: community features provide a robust meso-scale signal, while NF/LPF offer targeted gains when local neighborhood information is label-informative, without changing the ATLAS inference pipeline.
Convergence. ATLAS converges rapidly and stably across large graphs: training loss decreases smoothly, and validation performance plateaus early with a small train–validation gap. The curves exhibit no late-epoch degradation and remain stable after convergence (see Fig. 4).
Efficiency. Table 3 shows that ATLAS adds a $T$-resolution community search as a one-time preprocessing step with cost $O(T\|A\|_{0})$, after which training is MLP-like and inference is adjacency-free on features of dimension $D+Td_{c}$. Consequently, both preprocessing and inference costs increase with $T$. On OGBN-Products (Table 4), the larger $T$ leads to a noticeable preprocessing cost, yet per-epoch training remains competitive and inference stays sub-second. Appendix Table 5 reports timings on the remaining large graphs.
5 Accuracy Under Community Refinement
We quantify the level of community refinement by the minimum modularity threshold $Q_{\min}$. Large $Q_{\min}$ preserves only coarse communities; lowering $Q_{\min}$ progressively adds medium and fine communities, yielding a multi-scale representation. We define structural bias as how strongly a graph’s community structure provides a useful structural signal for classification, and group graphs into three regimes:
  • High structural bias (e.g., Cora, Tolokers): Community structure is strongly aligned with labels, so refinement helps. Coarse communities at large $Q_{\min}$ already carry substantial signal, and adding medium- and fine-grained communities reveals additional useful structure. As $Q_{\min}$ decreases and more refined communities are included, performance steadily improves until it saturates; see Figure 5 (top left and bottom left). In this regime, GCN and ATLAS both outperform MLP.
  • Low structural bias (e.g., Amazon-Ratings, Chameleon-Filtered, Flickr): Community structure is only weakly label-aligned. Coarse communities capture most of this limited structural signal, and adding finer communities yields at most small additional gains. As $Q_{\min}$ decreases and more refined communities are included, ATLAS improves moderately over the MLP, while GCN shows at best small or sometimes no gains over MLP; see Figure 5 (top center and bottom center).
  • Negative structural bias (e.g., Actor, Squirrel-Filtered, Roman-Empire): Community structure is misaligned with labels, and finer communities introduce noisy or misleading locality, so refinement hurts. As $Q_{\min}$ is lowered and more fine-grained communities are added, performance deteriorates; see Figure 5 (top right and bottom right). In this regime, GCN typically underperforms the MLP baseline.
Example (Cora). For Cora (Fig. 5 top left; Table 7), each choice of $Q_{\min}$ selects a subset of resolution parameters: we include community features from all resolutions whose modularity satisfies $Q(\gamma)\!\geq\!Q_{\min}$. When $Q_{\min}\!\in\!\{1.0,0.9\}$ no resolution meets the threshold, so ATLAS collapses to the feature-only MLP at $76.61\%$, below the GCN curve. At $Q_{\min}\!=\!0.8$, two resolutions are added and accuracy rises to $79.93\%$; at $0.7$ a medium-resolution setting increases it to $83.66\%$; and at $0.6$ two finer resolutions push it to $86.50\%$. As $Q_{\min}$ is lowered further and more resolutions are added, the ATLAS curve eventually overtakes GCN, reaching its peak of $88.10\%$ at $Q_{\min}\!=\!0.1$, where node features are augmented with a balanced mix of coarse, medium, and fine community features. Reducing $Q_{\min}$ to $0.0$ adds the most fragmented resolution and causes a slight drop, indicating diminishing returns from very fine community structure, which effectively acts as noise.
Figure 5: Effect of cumulatively adding community-derived features as the minimum modularity threshold $Q_{\min}$ is lowered, for high structural bias graphs (left), low structural bias graphs (middle), and negative structural bias graphs (right).
(a) Cora
(b) Tolokers
(c) Amazon-Ratings
(d) Chameleon-Filtered
(e) Actor
(f) Roman-Empire
Figure 6: NMI, mutual information, and entropy dynamics across resolutions. high structural bias datasets (Cora and Tolokers, left); low structural bias datasets (Amazon and Chameleon-Filtered, middle); negative structural bias datasets (Actor and Roman-Empire, right).
Figure 6 illustrates the refinement behavior predicted by our NMI theory. As resolution $\gamma$ increases, communities are refined and, as Lemmas 1–2 state, both $I(L;C)$ and $H(C)$ grow monotonically across all datasets. In contrast, Theorem 1 explains why NMI behaves differently across structural-bias regimes: on high structural-bias graphs it forms a clear interior peak, on low structural-bias graphs it stays low and fairly flat because gains in $I(L;C)$ do not outpace the entropy increase, and on negative structural-bias graphs it rises slowly from near zero and saturates at a modest level.
6 Ablation study
Effect of Modularity Gap and Resolution on Performance. The behavior of refinement levels––how many Louvain resolution values are selected and how useful they are––is strongly shaped by the modularity gap. A small average modularity gap keeps many closely spaced resolutions, while a large gap leaves only a few widely separated ones. This creates a trade-off between having many redundant community partitions and having too few, overly coarse partitions. On high structural-bias graphs (Figure 7(a)), accuracy is lowest at the extremes of this trade-off. When the average gap is very small, ATLAS retains many nearly redundant partitions and the additional community features mostly inject noise, depressing accuracy.As the gap moves into a moderate region ($gap\approx 0.06{-}0.09$), the selected resolutions capture community structure at several distinct granularities and align better with the labels, so mutual information strengthens and accuracy improves. If the gap becomes too large ($gap\gtrsim 0.10$), only a handful of coarse resolutions remain; the community structure is too crude to fully exploit the available signal and performance falls again.
For low structural-bias graphs (Figure 7(b)), the accuracy curves are much flatter. In this setting, the community structure carries little information about the labels, so changing the modularity gap mostly just changes how many community resolutions are kept, without making them much more predictive. As a result, adding community features yields only modest gains over a feature-only MLP, and accuracy is only weakly affected by the choice of gap.
Figure 7(c) illustrates this behavior on Cora. For small gaps, many closely spaced resolutions are selected and their communities are highly overlapping, so the extra features are largely redundant and behave as noise, matching the low-accuracy regime in Figure 7(a). For large gaps, only a few coarse resolutions remain and the community information is too crude to capture label-relevant structure. The best performance occurs at intermediate gaps, where a small set of resolutions captures different levels of granularity, providing informative structural signal without redundancy.
Figure 7: (a–b) Accuracy/ROC–AUC vs. average modularity gap on high- and low-structural-bias datasets. (c) Resolution count vs. modularity gap for Cora.
7 Conclusion and Future Work
We presented ATLAS, a community-augmented learning framework that enriches node features with multi-resolution Louvain embeddings and trains a compact MLP classifier. An adaptive resolution search, governed by $Q_{\min}$ and $\Delta Q$, selects a small set of informative resolutions, balancing coverage with cost. Across Q1 (homophily-regime benchmarks) and Q2 (large graphs), ATLAS attains competitive or superior accuracy relative to homophilic GNNs and heterophily-oriented models, while exhibiting fast, stable convergence and a favorable training footprint once preprocessing is complete.
In the future we aim to reduce preprocessing by improving resolution selection. A complementary direction is community-guided graph rewiring: using the discovered communities to propose label-aware edge edits that amplify useful intra-/inter-community signals and further improve accuracy.
Reproducibility Statement. Our source code is available at https://github.com/atlaspaper16/ATLAS.